OpenStudy (anonymous):

in triangle abc bd perpendicular to ac,ac^2-bc^2=bc^2.prove that bd^2=ad*dc

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

i think you have written the statement wrong.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\left( ac \right)^{2}-\left( ab \right)^{2}=\left( bc \right)^{2}\]